{\small
\begin{thebibliography}{10}

\bibitem{alzahrani2016spatio}
N.~Alzahrani, M.~Spichkova, and J.~O. Blech.
\newblock Spatio-temporal models for formal analysis and property-based
  testing.
\newblock In {\em Federation of International Conferences on Software
  Technologies: Applications and Foundations}, pages 196--206. Springer, 2016.

\bibitem{alzahrani2017temporal}
N.~Alzahrani, M.~Spichkova, and J.~O. Blech.
\newblock From temporal models to property-based testing.
\newblock In {\em 11th International Conference on Evaluation of Novel
  Approaches to Software Engineering (ENASE)}, pages 241--246. SCITEPRESS,
  2017.

\bibitem{aravantinos2015autofocus}
V.~Aravantinos, S.~Voss, S.~Teufl, F.~H{\"o}lzl, and B.~Sch{\"a}tz.
\newblock {AutoFOCUS 3: Tooling Concepts for Seamless, Model-based Development
  of Embedded Systems}.
\newblock In {\em ACES-MB}, pages 19--26, 2015.

\bibitem{bohm2014formal}
W.~B{\"o}hm, M.~Junker, A.~Vogelsang, S.~Teufl, R.~Pinger, and K.~Rahn.
\newblock A formal systems engineering approach in practice: An experience
  report.
\newblock In {\em Proceedings of the 1st International Workshop on Software
  Engineering Research and Industrial Practices}, pages 34--41. ACM, 2014.

\bibitem{botaschanjan2008correctness}
J.~Botaschanjan, M.~Broy, A.~Gruler, A.~Harhurin, S.~Knapp, L.~Kof, W.~Paul,
  and M.~Spichkova.
\newblock On the correctness of upper layers of automotive systems.
\newblock {\em Formal aspects of computing}, 20(6):637--662, 2008.

\bibitem{botaschanjan2006towards}
J.~Botaschanjan, A.~Gruler, A.~Harhurin, L.~Kof, M.~Spichkova, and
  D.~Trachtenherz.
\newblock Towards modularized verification of distributed time-triggered
  systems.
\newblock In {\em Int. Symposium on Formal Methods}, pages 163--178. Springer,
  2006.

\bibitem{botaschanjan2005towards}
J.~Botaschanjan, L.~Kof, C.~K{\"u}hnel, and M.~Spichkova.
\newblock Towards verified automotive software.
\newblock In {\em ACM SIGSOFT Software Engineering Notes}, volume~30, pages
  1--6. ACM, 2005.

\bibitem{brock1981scenarios}
J.~D. Brock and W.~B. Ackerman.
\newblock Scenarios: A model of non-determinate computation.
\newblock In {\em Formalization of programming concepts}, pages 252--259.
  Springer, 1981.

\bibitem{broy1995advanced}
M.~Broy.
\newblock Advanced component interface specification.
\newblock {\em Theory and Practice of Parallel Programming}, pages 369--392,
  1995.

\bibitem{broy1995buffer}
M.~Broy.
\newblock Specification and refinement of a buffer of length one.
\newblock In {\em Deductive program design. NATO ASI Series, Series F: Computer
  and Systems Sciences. Vol. 152}. Springer, 1995.

\bibitem{broy_refinement}
M.~Broy.
\newblock Compositional refinement of interactive systems.
\newblock {\em ACM}, 44(6):850--891, 1997.

\bibitem{broy1997component}
M.~Broy.
\newblock Towards a mathematical concept of a component and its use.
\newblock {\em Software-concepts and tools}, 18(3):137, 1997.

\bibitem{Broy99alogical}
M.~Broy.
\newblock A logical basis for component-based systems engineering.
\newblock In {\em Calculational System Design. IOS}. Press, 1999.

\bibitem{broy_time_abstr}
M.~Broy.
\newblock {Time, Abstraction, Causality and Modularity in Interactive Systems:
  Extended Abstract}.
\newblock {\em Electr. Notes Theor. Comput. Sci.}, 108:3--9, 2004.

\bibitem{broy_janus}
M.~Broy.
\newblock {Service-oriented Systems Engineering: Specification and design of
  services and layered architectures. {T}he {JANUS} {A}pproach}.
\newblock {\em Engineering Theories of Software Intensive Systems}, pages
  47--81, 2005.

\bibitem{Broy2010MSS}
M.~Broy.
\newblock Structured modeling and specification of functional requirements.
\newblock {\em Science of Computer Programming}, 75(12):1193--1214, 2010.

\bibitem{broy1997summary}
M.~Broy, M.~Breitling, B.~Sch{\"a}tz, and K.~Spies.
\newblock Summary of case studies in focus: A design method for distributed
  systems.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, I9740,
  1997.

\bibitem{broy1992summary}
M.~Broy, F.~Dederichs, C.~Dendorfer, M.~Fuchs, T.~Gritzner, and R.~Weber.
\newblock Summary of case studies in focus: A design method for distributed
  systems.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 9203, 1992.

\bibitem{broy1992design}
M.~Broy, F.~Dederichs, C.~Dendorfer, M.~Fuchs, T.~F. Gritzner, and R.~Weber.
\newblock {The design of distributed systems: an introduction to Focus}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 9202, 1992.

\bibitem{cocome}
M.~Broy, J.~Fox, F.~H\"{o}lzl, D.~Koss, M.~Kuhrmann, M.~Meisinger,
  B.~Penzenstadler, S.~Rittmann, B.~Sch\"{a}tz, M.~Spichkova, and D.~Wild.
\newblock {Service-Oriented Modeling of CoCoME with Focus and AutoFocus}.
\newblock pages 177--206, 2008.

\bibitem{broy1994summary}
M.~Broy, M.~Fuchs, T.~Gritzner, B.~Sch{\"a}tz, and K.~S. Katharina~Spies.
\newblock Summary of case studies in focus: A design method for distributed
  systems.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 9423, 1994.

\bibitem{broy1999autof}
M.~Broy, F.~Huber, and B.~Sch{\"a}tz.
\newblock {AutoFocus -- Ein Werkzeugprototyp zur Entwicklung eingebetteter
  Systeme}.
\newblock {\em Informatik-Forschung und Entwicklung}, 14(3):121--134, 1999.

\bibitem{Broy2007services}
M.~Broy, I.~H. Kr\"{u}ger, and M.~Meisinger.
\newblock A formal model of services.
\newblock {\em ACM Trans. Softw. Eng. Methodol.}, 16(1), Feb. 2007.

\bibitem{broy1999boiler}
M.~Broy, F.~Regensburger, B.~Sch{\"a}tz, and K.~Spies.
\newblock {The Steamboiler Specification -- A Case Study in Focus}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 9202, 1992.

\bibitem{broy1994specification}
M.~Broy and K.~St{\o}len.
\newblock Specification and refinement of finite dataflow networks -- a
  relational approach.
\newblock In {\em Formal Techniques in Real-Time and Fault-Tolerant Systems},
  pages 247--267. Springer, 1994.

\bibitem{focus}
M.~Broy and K.~St{\o}len.
\newblock {\em Specification and Development of Interactive Systems: Focus on
  Streams, Interfaces, and Refinement}.
\newblock Springer, 2001.

\bibitem{campetelli2011user}
A.~Campetelli, F.~H{\"o}lzl, and P.~Neubeck.
\newblock User-friendly model checking integration in model-based development.
\newblock {\em The International Society for Computers and Their Applications},
  2011.

\bibitem{campetelli2015model}
A.~Campetelli, M.~Junker, B.~B{\"o}hm, M.~Davidich, V.~Koutsoumpas, X.~Zhu, and
  J.~C. Wehrstedt.
\newblock A model-based approach to formal verification in early development
  phases: A desalination plant case study.
\newblock In {\em Software Engineering}, pages 91--100, 2015.

\bibitem{cimatti2016temporal}
A.~Cimatti and S.~Tonetta.
\newblock A temporal logics approach to contract-based design.
\newblock In {\em Architecture-Centric Virtual Integration (ACVI)}, pages 1--3.
  IEEE, 2016.

\bibitem{dobi2015model}
S.~Dobi, M.~Gleirscher, M.~Spichkova, and P.~Struss.
\newblock Model-based hazard and impact analysis.
\newblock {\em arXiv preprint arXiv:1512.02759}, 2015.

\bibitem{feilkas2009top}
M.~Feilkas, A.~Fleischmann, F.~H{\"o}lzl, C.~Pfaller, K.~Scheidemann,
  M.~Spichkova, and D.~Trachtenherz.
\newblock A top-down methodology for the development of automotive software.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 902, 2009.

\bibitem{feilkas2011refined}
M.~Feilkas, F.~H{\"o}lzl, C.~Pfaller, S.~Rittmann, B.~Sch{\"a}tz, W.~Schwitzer,
  W.~Sitou, M.~Spichkova, and D.~Trachtenherz.
\newblock {A Refined Top-Down Methodology for the Development of Automotive
  Software Systems -- The KeylessEntry System Case Study}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 1103, 2011.

\bibitem{holzl201013}
F.~H{\"o}lzl and M.~Feilkas.
\newblock Autofocus 3 -- a scientific tool prototype for model-based
  development of component-based, reactive, distributed systems.
\newblock In {\em Model-Based Engineering of Embedded Real-Time Systems}, pages
  317--322. Springer, 2010.

\bibitem{holzl2010autofocus}
F.~H{\"o}lzl, M.~Spichkova, and D.~Trachtenherz.
\newblock {AutoFocus Tool Chain}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.},
  (TUM-I1021), 2010.

\bibitem{holzl2010safety}
F.~H{\"o}lzl, M.~Spichkova, and D.~Trachtenherz.
\newblock Safety-critical system development methodology.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, (1020),
  2010.

\bibitem{AF3modevva2017}
S.~Kanav and V.~Aravantinos.
\newblock Modular transformation from af3 to nuxmv.
\newblock In {\em MoDeVVa 2017}. CEUR, 2017.

\bibitem{kondeva2013seamless}
A.~Kondeva, D.~Ratiu, B.~Sch{\"a}tz, and S.~Voss.
\newblock {Seamless model-based development of embedded systems with AF3
  phoenix}.
\newblock In {\em Engineering of Computer Based Systems (ECBS)}, pages
  212--212. IEEE, 2013.

\bibitem{kuhnel2006flexray}
C.~K{\"u}hnel and M.~Spichkova.
\newblock {FlexRay und FTCom: Formale Spezifikation in FOCUS}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep. I}, 601:2006,
  2006.

\bibitem{kuhnel2006upcoming}
C.~K{\"u}hnel and M.~Spichkova.
\newblock {Upcoming automotive standards for fault-tolerant communication:
  FlexRay and OSEKtime FTCom}.
\newblock In {\em EFTS 2006 International Workshop on Engineering of Fault
  Tolerant Systems. Universite du Luxembourg, CSC: Computer Science and
  Communication}, 2006.

\bibitem{kuhnel2007fault}
C.~K{\"u}hnel and M.~Spichkova.
\newblock Fault-tolerant communication for distributed embedded systems.
\newblock In {\em Software Engineering of Fault Tolerance Systems (Series on
  Software Engineering and Knowledge Engineering)}, volume~19, page 175. World
  Scientific Publishing, 2007.

\bibitem{npw}
T.~Nipkow, L.~C. Paulson, and M.~Wenzel.
\newblock {\em {Isabelle/HOL -- A Proof Assistant for Higher-Order Logic}},
  volume 2283 of {\em LNCS}.
\newblock Springer, 2002.

\bibitem{philipps2014stepwise}
J.~Philipps and B.~Rumpe.
\newblock Stepwise refinement of data flow architectures.
\newblock {\em arXiv preprint arXiv:1409.7241}, 2014.

\bibitem{schatz1996graphical}
B.~Sch{\"a}tz, H.~Hu{\ss}mann, and M.~Broy.
\newblock Graphical development of consistent system specifications.
\newblock In {\em FME'96: Industrial Benefit and Advances In Formal Methods},
  pages 248--267. Springer, 1996.

\bibitem{simic2016enhancing}
M.~Simic, M.~Spichkova, H.~Schmidt, and I.~Peake.
\newblock Enhancing learning experience by collaborative industrial projects.
\newblock In {\em ICEER 2016}, pages 1--8. Western Sydney University, 2016.

\bibitem{spichkova2006flexray}
M.~Spichkova.
\newblock {FlexRay: Verification of the FOCUS Specification in Isabelle/HOL. A
  Case Study}.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, (602),
  2006.

\bibitem{spichkova}
M.~Spichkova.
\newblock {\em {Specification and Seamless Verification of Embedded Real-Time
  Systems: FOCUS on Isabelle}}.
\newblock PhD thesis, TU M{\"u}nchen, 2007.

\bibitem{spichkova2008focus}
M.~Spichkova.
\newblock Focus on isabelle: From specification to verification.
\newblock {\em Department of Electrical and Computer Engineering, Concordia
  University, Tech. Rep}, 2008.

\bibitem{spichkova2008refinement}
M.~Spichkova.
\newblock Refinement-based verification of interactive real-time systems.
\newblock {\em Electronic Notes in Theoretical Computer Science}, 214:131--157,
  2008.

\bibitem{spichkova2009refinement}
M.~Spichkova.
\newblock Refinement-based specification and verification (short version).
\newblock In {\em Dagstuh Research Seminar on Refinement Based Methods for the
  Construction of Dependable Systems}, 2009.

\bibitem{spichkova2010architecture}
M.~Spichkova.
\newblock Architecture: Methodology of decomposition.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.},
  (TUM-I1018), 2010.

\bibitem{spichkova2010semiformal}
M.~Spichkova.
\newblock {From Semiformal Requirements To Formal Specifications via MSCs}.
\newblock Technical report, Technische Universit{\"a}t M{\"u}nchen, 2010.

\bibitem{spichkova2011decomp}
M.~Spichkova.
\newblock {Architecture: Requirements + Decomposition + Refinement}.
\newblock {\em Softwaretechnik-Trends}, 31:4, 2011.

\bibitem{spichkova2011focus}
M.~Spichkova.
\newblock Focus on processes.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.},
  (TUM-I1115), 2011.

\bibitem{spichkova2011refinement}
M.~Spichkova.
\newblock Refinement-based specification: Requirements and architecture.
\newblock In {\em Software Engineering 2011 (SE)}. Karlsruher Institut f{\"u}r
  Technologie, 2011.

\bibitem{spichkova2012component}
M.~Spichkova.
\newblock Component composition: Formal specification and verification of
  cryptographic properties.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, (TUM-I124),
  2012.

\bibitem{hffm_spichkova}
M.~Spichkova.
\newblock {Human Factors of Formal Methods}.
\newblock In {\em {IADIS Interfaces and Human Computer Interaction}}. IHCI,
  2012.

\bibitem{spichkova2012time}
M.~Spichkova.
\newblock {Towards Focus on Time}.
\newblock In {\em {12th International Workshop on Automated Verification of
  Critical Systems (AVoCS)}}, 2012.

\bibitem{spichkova2013design}
M.~Spichkova.
\newblock Design of formal languages and interfaces: “formal” does not mean
  “unreadable”.
\newblock In {\em Emerging Research and Trends in Interactivity and the
  Human-Computer Interface}. IGI Global, 2013.

\bibitem{FocusStreamsCaseStudies-AFP}
M.~Spichkova.
\newblock {Stream Processing Components: Isabelle/HOL Formalisation and Case
  Studies}.
\newblock {\em {Archive of Formal Proofs}}, 2013.

\bibitem{CryptoBasedCompositionalProperties-AFP}
M.~Spichkova.
\newblock Compositional properties of crypto-based components.
\newblock {\em Archive of Formal Proofs}, 2014.

\bibitem{spichkova2014formalisation}
M.~Spichkova.
\newblock Formalisation and analysis of component dependencies.
\newblock {\em Archive of Formal Proofs}, 2014.

\bibitem{spichkova2015trade}
M.~Spichkova.
\newblock Trade: Formal modelling and analysis framework.
\newblock In {\em 4th International Symposium on Modelling and Knowledge},
  2015.

\bibitem{spichkova2016spatio}
M.~Spichkova.
\newblock {Spatio-temporal features of Focus$^{ST}$}.
\newblock {\em arXiv preprint arXiv:1610.07884}, 2016.

\bibitem{spichkova2014modeling}
M.~Spichkova, J.~Blech, P.~Herrmann, and H.~Schmidt.
\newblock {Modeling spatial aspects of safety-critical systems with
  Focus$^{ST}$}.
\newblock In {\em MoDeVVa2014}, pages 49--58. CEUR, 2014.

\bibitem{spichkova2012cyber}
M.~Spichkova and A.~Campetelli.
\newblock Towards system development methodologies: From software to
  cyber-physical domain.
\newblock In {\em First International Workshop on Formal Techniques for
  Safety-Critical Systems (FTSCS'12)}, 2012.

\bibitem{spichkova2012verified}
M.~Spichkova, F.~H{\"o}lzl, and D.~Trachtenherz.
\newblock Verified system development with the autofocus tool chain.
\newblock In {\em 2nd Workshop on Formal Methods in the Development of Software
  (WS-FMDS 2012)}, volume~86, pages 17--24. Electronic Proceedings in
  Theoretical Computer Science, 2012.

\bibitem{spichkova2008formal}
M.~Spichkova and J.~J{\"u}rjens.
\newblock Formal specification of cryptographic protocols and their composition
  properties: Focus-oriented approach.
\newblock Technical report, Technische Universit{\"a}t M{\"u}nchen, 2008.

\bibitem{spichkova2015human}
M.~Spichkova, H.~Liu, M.~Laali, and H.~W. Schmidt.
\newblock Human factors in software reliability engineering.
\newblock In {\em Workshop on Applications of Human Error Research to Improve
  Software Engineering (WAHESE2015)}, 2015.

\bibitem{spichkova2014towards}
M.~Spichkova and H.~Schmidt.
\newblock Towards logical architecture and formal analysis of dependencies
  between services.
\newblock In {\em The 2014 Asia-Pacific Services Computing Conference}, 2014.

\bibitem{spichkova2015formal}
M.~Spichkova and H.~Schmidt.
\newblock Formal-based framework for analysis of logical architecture.
\newblock {\em International Journal of Services Computing (IJSC)}, 3(3):1--15,
  2015.

\bibitem{mise2015process}
M.~Spichkova and H.~Schmidt.
\newblock Reconciling a component and process view.
\newblock {\em 7th International Workshop on Modeling in Software Engineering
  (MiSE) at ICSE 2015}, 2015.

\bibitem{spichkova2016managing}
M.~Spichkova, H.~Schmidt, I.~E. Thomas, I.~I. Yusuf, S.~G. Androulakis, and
  G.~R. Meyer.
\newblock Managing usability and reliability aspects in cloud computing.
\newblock In {\em 11th International Conference on Evaluation of Novel
  Approaches to Software Engineering}, pages 288--295, 2016.

\bibitem{spichkova2017towards}
M.~Spichkova, H.~W. Schmidt, I.~I. Yusuf, I.~E. Thomas, S.~Androulakis, and
  G.~R. Meyer.
\newblock Towards modelling and implementation of reliability and usability
  features for research-oriented cloud computing platforms.
\newblock In {\em Evaluation of Novel Approaches to Software Engineering.
  Series on Communications in Computer and Information Science}, pages
  158--178. Springer, 2017.

\bibitem{spichkova2015towards}
M.~Spichkova and M.~Simic.
\newblock Towards formal modelling of autonomous systems.
\newblock In {\em Intelligent Interactive Multimedia Systems and Services},
  pages 279--288. Springer, 2015.

\bibitem{spichkova2017autonomous}
M.~Spichkova and M.~Simic.
\newblock Autonomous systems research embedded in teaching.
\newblock In {\em Intelligent Interactive Multimedia Systems and Services},
  pages 268--277. Springer, 2017.

\bibitem{spichkova2017human}
M.~Spichkova and M.~Simic.
\newblock Human-centred analysis of the dependencies within sets of proofs.
\newblock In {\em Knowledge-Based and Intelligent Information \& Engineering
  Systems}, pages 2290--2298. Elsevier, 2017.

\bibitem{spichkova2015route}
M.~Spichkova, M.~Simic, and H.~Schmidt.
\newblock Formal model for intelligent route planning.
\newblock {\em Procedia Computer Science}, 60:1299--1308, 2015.

\bibitem{spichkova2016automotive}
M.~Spichkova, M.~Simic, and H.~Schmidt.
\newblock From automotive to autonomous: Time-triggered operating systems.
\newblock In {\em Intelligent Interactive Multimedia Systems and Services
  2016}, pages 347--359. Springer, 2016.

\bibitem{spichkova2016formal}
M.~Spichkova, M.~Simic, H.~Schmidt, J.~Cheng, X.~Dong, Y.~Gui, Y.~Liang,
  P.~Ling, and Z.~Yin.
\newblock Formal models for intelligent speed validation and adaptation.
\newblock {\em Procedia Computer Science}, 96:1609--1618, 2016.

\bibitem{ChimineyICPADS}
M.~Spichkova, I.~Thomas, H.~Schmidt, I.~Yusuf, D.~Drumm, S.~Androulakis,
  G.~Opletal, and S.~Russo.
\newblock Scalable and fault-tolerant cloud computations: Modelling and
  implementation.
\newblock In {\em 21st IEEE International Conference on Parallel and
  Distributed Systems}, 2015.

\bibitem{spichkova2016teaching}
M.~Spichkova and A.~Zamansky.
\newblock Teaching formal methods for software engineering.
\newblock In {\em ENASE 2016}, pages 370--376. Science and Technology
  Publications, 2016.

\bibitem{spichkova2013we}
M.~Spichkova, X.~Zhu, and D.~Mou.
\newblock Do we really need to write documentation for a system?
\newblock In {\em International Conference on Model-Driven Engineering and
  Software Development (MODELSWARD'13)}, 2013.

\bibitem{Spivey_88}
M.~Spivey.
\newblock {Understanding Z -- A Specification Language and Its Formal
  Semantics}.
\newblock Cambridge Tracts in Theoretical Computer Science 3. Camb. Univ.
  Press, 1988.

\bibitem{Spivey_92}
M.~Spivey.
\newblock {The Z Notation: A Reference Manual, 2. Ausgabe}.
\newblock Prentice-Hall International Series in Computer Science.
  Prentice-Hall, 1992.

\bibitem{stolen1996using}
K.~St{\o}len.
\newblock Using relations on streams to solve the rpc-memory specification
  problem.
\newblock {\em Formal Systems Specification}, pages 477--520, 1996.

\bibitem{teufl2013mira}
S.~Teufl, D.~Mou, and D.~Ratiu.
\newblock Mira: A tooling-framework to experiment with model-based requirements
  engineering.
\newblock In {\em Requirements Engineering Conference (RE), 2013 21st IEEE
  International}, pages 330--331. IEEE, 2013.

\bibitem{vo2016model}
P.~Vo and M.~Spichkova.
\newblock Model-based generation of natural language specifications.
\newblock {\em Software Technologies: Applications and Foundations},
  (Human-Oriented Formal Methods), 2016.

\bibitem{vogelsang2014supporting}
A.~Vogelsang, S.~Eder, G.~Hackenberg, M.~Junker, and S.~Teufl.
\newblock Supporting concurrent development of requirements and architecture: A
  model-based approach.
\newblock In {\em Model-Driven Engineering and Software Development}, pages
  587--595. IEEE, 2014.

\bibitem{voss2012scheduling}
S.~Voss and B.~Sch{\"a}tz.
\newblock {Scheduling shared memory multicore architectures in AutoFocus 3
  using Satisfiability Modulo Theories}.
\newblock In {\em Modellbasierte Entwicklung eingebetteter Systeme}, page~49,
  2012.

\bibitem{voss2013deployment}
S.~Voss and B.~Sch{\"a}tz.
\newblock Deployment and scheduling synthesis for mixed-critical shared-memory
  applications.
\newblock In {\em Engineering of Computer Based Systems}, pages 100--109. IEEE,
  2013.

\bibitem{walicki1995structured}
M.~Walicki and M.~Broy.
\newblock Structured specifications and implementation of nondeterministic data
  types.
\newblock {\em Technische Universit{\"a}t M{\"u}nchen, Tech. Rep.}, 9442, 1995.

\bibitem{Isar}
M.~Wenzel.
\newblock {The Isabelle/Isar Reference Manual}.
\newblock http://isabelle.in.tum.de/doc/isar-ref.pdf, 2016.

\bibitem{Chiminey}
I.~Yusuf, I.~Thomas, M.~Spichkova, S.~Androulakis, G.~Meyer, D.~Drumm,
  G.~Opletal, S.~Russo, A.~Buckle, and H.~Schmidt.
\newblock Chiminey: Reliable computing and data management platform in the
  cloud.
\newblock In {\em International Conference on Software engineering}, pages
  677--680, 2015.

\bibitem{yusuf2017chiminey}
I.~I. Yusuf, I.~E. Thomas, M.~Spichkova, and H.~W. Schmidt.
\newblock Chiminey: Connecting scientists to hpc, cloud and big data.
\newblock {\em Big Data Research}, 8:39--49, 2017.

\bibitem{zamansky2016formal}
A.~Zamansky, G.~Rodriguez-Navas, M.~Adams, and M.~Spichkova.
\newblock Formal methods in collaborative projects.
\newblock In {\em 11th International Conference on Evaluation of Novel
  Approaches to Software Engineering. IEEE}, 2016.

\end{thebibliography}
}

\documentclass[a4paper,twoside]{article}

\usepackage{epsfig}
\usepackage{subfigure}
\usepackage{calc}
\usepackage{amssymb}
\usepackage{amstext} 
\usepackage{amsthm} 
\usepackage{xspace}
\usepackage{focus}
\usepackage[table,xcdraw]{xcolor}  
\usepackage{textcomp}
\usepackage{url}
\usepackage{array,multirow,graphicx}

\usepackage{SCITEPRESS}      

\begin{document}

\newcommand{\focust}{\textsc{Focus}$^{ST}$}
\newcommand{\focuse}{\textsc{Focus}$^{E}$}

\title{(Auto)Focus approaches and their applications:\\
A systematic review}

\keywords{Software Engineering, Formal Methods, Specification, Verification, Tool-support}

\author{\authorname{Maria Spichkova}
\affiliation{School of Science, RMIT University, 414-418 Swanston Street, 3001, Melbourne, Australia} 
 \email{maria.spichkova@rmit.edu.au}
}

\abstract{
\Focus,
a framework for formal specification and development of interactive systems, 
was introduced approx. 25 years ago.
Since then this approach was broadly used in academic and industrial studies, as well as provided a basis  for a number of another frameworks focusing on particular domains, and  
for the AF3 modelling tool. 
In this paper we provide a literature review of the corresponding approaches, academic case studies and industrial applications of these methods.
}

\onecolumn 
\maketitle 
\normalsize \vfill

\section{\uppercase{Introduction}}

\Focus,
a framework for formal specification and development of distributed interactive systems, 
provides models and formalisms for a step-wise specification and development. 
\Focus was introduced by Broy et al. \cite{broy1992design} approx. 25 years ago. Since then, the framework was extended and gave a basis for a number of other approaches. Many case studies from several application domains were conducted, within both academic and industrial projects. 
One of the most comprehensive description of the formal background of the framework was presented by Broy and St{\o}len \cite{focus}, 
and became a highly cited publication (more than 540 citations\footnote{According to the Google Scholar, retr. 20/11/17}). 
The goal of this approach is to support a modular system development with the help of component specifications, having precisely identified interfaces and formal refinement concepts. 
Focus is based on  a clear mathematical theory and aims on system development carried out in a systematic way. 

The tool-support for this approach is provided by the AutoFocus CASE tool.
The prototype was  initially presented by Broy et al. \cite{broy1999autof}, and later extended and refined by two next generations of the tool. 
The third  generation of AutoFocus, AutoFocus3 or AF3,
is a powerful open-source  tool.  
It's emphasis is on the development of embedded systems using models from the requirements to the hardware architecture. 
The latest version of AF3 provides advanced features to support the user: 
formal analyses, synthesis methods, space exploration visualization, etc.

The research on the \Focus-related approaches is an active topic. 
The most resent related works were presented by Alzahrani et al.
\cite{alzahrani2017temporal}, where aims was to apply property-based testing on formal models with temporal properties, and by Kanav and Aravantinos
\cite{AF3modevva2017}, introducing the modular transformation from AF3 to the \emph{nuXmv} symbolic model checker.
As over the last 25 years many theoretical and applies work was conducted within the \Focus approach, an overview of existing work is required.  
In this paper, we are going to provide a systematic review of the corresponding approaches as well as on the case studies they were applied on.

\emph{Outline:} 
Section \ref{sec:focus} provided a high-level overview of the \Focus ideas. 
Spatio-Temporal view on modelling of interactive systems is presented in Section \ref{sec:focust}.
Section \ref{sec:auto} discusses the tool support for the approach.
Section \ref{sec:methodologies} introduces methodological extensions 
of the \Focus approach as well the frameworks built on its basis.  
Section \ref{sec:applications} reviews the academic and industrial case studies conducted using \Focus or related approaches.
Section \ref{sec:hofm} discusses human-oriented aspects of the related formal methods.
Finally, Section \ref{sec:conclusions} summarises the paper.

\section{\uppercase{Focus}}
\label{sec:focus}
 
A distributed system in \Focus is specified by its logical \emph{components} 
 connected via \emph{channels}, which  are  directed and order preserving. 
Components  work independently of each other or interact, exchanging
  information in terms of \emph{messages} of specified types.
  
The formal meaning of a \Focus specification is an \emph{input/output relation}, i.e. 
 the relation between the
communication histories for the external \emph{input} and \emph{output channels}. 
The specifications can be 
\begin{itemize}
\item structured into a number of
formulas each characterizing a different kind of property;
\item elementary or composite: 
  \begin{itemize}
  \item Elementary specifications can be
untimed, timed, and time-synchronous;
  \item Composite specifications  are
built hierarchically from the elementary ones.
  \end{itemize}
\end{itemize}
The central concept in \Focus are \emph{streams}, that represent 
communication histories of \emph{directed channels}. 
Streams in \Focus are functions mapping the indexes in their domains
 to their messages. 
For any set of messages $M$, $M^\omega$ denotes the set of all streams,
$M^\infty$ and $M^*$ denote  the sets of all infinite and all finite
streams respectively.
 $M^{\underline{\omega}}$ denotes the set of all 
timed streams, $M^{\underline{\infty}}$ and $M^{\underline{*}}$ denote the sets of 
all infinite and all finite timed streams respectively.

\[ \begin{array}{l}
M^\omega \stackrel{\mathrm{def}}{=} M^* \cup M^\infty 
\\
M^* \stackrel{\mathrm{def}}{=} {\bigcup}_{n \in \mathbb{N}}([1..n]\to M)
\\
M^\infty \stackrel{\mathrm{def}}{=} \mathbb{N}_{+} \to M 
\end{array}
\]
 Thus,  \emph{timed stream} is represented by a sequence of messages and {\it time ticks}, 
the messages are also listed in their order of transmission. The ticks model a discrete notion of time.  
\[ 
\begin{array}{l}
M^{\underline{\omega}}$ =  $M^{\underline{*}} \cup M^{\underline{\infty}}\\
M^{\underline{*}} \stackrel{\mathrm{def}}{=} 
\bigcup_{n \in \mathbb{N}}([1..n]\to M \cup \{ \surd \})\\
M^{\underline{\infty}} \stackrel{\mathrm{def}}{=} 
\mathbb{N}_+ \to M \cup \{ \surd \}
\end{array}
\] 

\section{\uppercase{Spatio-Temporal View: Focus$^{ST}$}}
\label{sec:focust}

Timing aspects of \Focus as well as the corresponding optimisations of the specification layout were discussed in \cite{spichkova2012time}, which was the first step towards elaboration the \focust framework.

In both frameworks, specifications are based on the notion of \emph{streams}.
However, in the original \Focus  input and output streams of a component are mappings
 of natural numbers $\Nat$ to single messages,
whereas a \focust\  stream %
 is a mapping from $\Nat$ to lists of messages  within the corresponding time intervals.
 Moreover, the syntax of \focust\ is particularly devoted to specify spatial (S) and timing (T) aspects in a comprehensible fashion,
which is the reason to extend the name of the language by $^{ST}$.

The \focust\ specification layout \cite{spichkova2016spatio,spichkova2014modeling} is similar to \Focus (which layout was inspired by Z specification language, cf. \cite{Spivey_88,Spivey_92}), 
but it has many new features to increase the readability and understandability of the specification.  
The \focust\ specification layout 
is based on human factor analysis within formal methods (see Section~\ref{sec:hofm}). 

In \focust specifications, input and output streams of a component are always timed, as spatio-temporal aspects are the core of the framework. 
The (timed) streams are mappings from $\Nat$ to lists of messages  within the corresponding time intervals. 
Thus, these streams are infinite per default, but they could be empty completely or from a certain point which is represented by empty time intervals $\nempty$.  
More precisely, \focust has streams of two kinds:
	\begin{itemize}
	\item
	\emph{Infinite timed streams} (denoted by $M^{\underline{\infty}}$) are used to represent the input and the output streams;
	\item
	\emph{finite timed streams} (denoted by $M^{\underline{*}}$) are used  
	to argue about a timed stream that was truncated 
	at some point of time.
	\end{itemize}  
 Infinite timed streams of type $T$ are defined  by a functional type 
\[
\Nat\ \to \nfst{T}
\] 
Finite timed streams of type  $T$  are defined by list of lists over this type, i.e., 
\[
\nfst{(\nfst{T})}
\]
where $\nfst{T}$ denotes a list of elements of type $T$.

The \focust ideas were applied within the approach on an intelligent route planning for a public transport system within a sustainable Smart City, cf.  \cite{spichkova2015trade,spichkova2015route}.
 
Spatio-temporal models for formal analysis and property-based testing were presented in 
\cite{alzahrani2016spatio,alzahrani2017temporal} by Alzahrani et al.
The authors aimed to to apply property-based testing on \focust and TLA models  with temporal properties.

\section{\uppercase{AutoFocus}}
\label{sec:auto}

AutoFocus tool was developed based on the \Focus theory. 
The first two prototype versions were replaced by AF3 (AutoFocus 3, cf. \cite{holzl201013}) tool that provides a various functionalities and supports several development phases.\footnote{\url{https://af3.fortiss.org/}} 
AF3 embeds the core modelling artefacts, is open source, and has a well defined formal syntax behind all its modelling elements.
Source code of AutoFocus3 models are coded in XML, which makes it easy to parse and to analyse.   
The tool was applied as a part of tool chain within a number of development methodologies, cf. Section  \ref{sec:methodologies}.

H{\"o}lzl et al. proposed an idea of AutoFocus tool chain and its application for the development of safety-critical systems \cite{holzl2010autofocus,holzl2010safety}. 
An extended and refined version of the approach was later presented in  \cite{spichkova2012verified}. 
This approach  also was elaborated using AutoFocus 2 and allows a direct application within AF3.
 
Campetelli et al. \cite{campetelli2011user} introduced an approach to a user-friendly model
checking integration in model-based development using AF3. 
The approach supports support
two different model checkers for the model and
implementation code verification: SMV and TVARC.

Kondeva et al. 
\cite{kondeva2013seamless} presented how the integrated system views on several levels of abstraction are implemented in AF3, to allow the development of embedded systems.

Teufl et al. \cite{teufl2013mira} presented an integrated with AF3 tooling-framework to experiment with model-based Requirements engineering.

Cimatti  and Tonetta \cite{cimatti2016temporal} introduced 
a temporal logics approach to contract-based design, integrated within AF3. 

AF3 tooling concepts for  
model-based development of embedded systems was presented in~\cite{aravantinos2015autofocus}.

A number of approaches on scheduling and deployment were also integrated within  AF3.
Voss and Sch{\"a}tz  presented an approach on scheduling shared memory multicore architectures in AF3 using SMT solvers \cite{voss2012scheduling}, as well as an approach
on deployment and scheduling Synthesis for mixed-critical shared-memory applications \cite{voss2013deployment}.

An approach on generating formal specifications from system AF3 models was presented in \cite{spichkova2013we}. 
Applying this approach would allow to solve  the problem with outdated system documentation  by making the documentation updates automatically: 
an up-to-date formal specification could be generated from the model if the model is frequently changed.  
The next step in this detection was presented in \cite{vo2016model}: this approach  allowed  generation of natural language specifications from AF3 models.

\section{\uppercase{Methodologies}}
\label{sec:methodologies}

A number of software and system development methodologies were introduced involving   \Focus or the framework created in its basis, as well as involving  an extensive use of AutoFocus tools.

One of the core features of the current \Focus framework, the specification of the black box behaviour of data flow components by characterizing the
relation between the input and the output histories, was initially introduced in \cite{broy1994specification}. The authors distinguished between three
main specification classes: time-independent specifications, weakly
time-dependent and strongly time-dependent specifications. 
Data flow components were formally specified by sets of timed stream processing functions.
Specifications describe such sets by logical formulas. The proposed solution allowed to handle the Brock/Ackermann anomaly~\cite{brock1981scenarios}.  
A further analysis of considering a component as a black box that
is a physical encapsulation of related services was presented in \cite{broy1997component}.
Philipps and Rumpe proposed a stepwise refinement of data flow architectures, cf. \cite{philipps2014stepwise}.

Another core feature of the current \Focus framework,  a method for the specification of reactive asynchronous components with a concurrent access interface, was discussed in
\cite{broy1995advanced}.

The approach for structured specifications and implementation of non-deterministic data types
was introduced in \cite{walicki1995structured}.   

St{\o}len used relations on \Focus streams to solve the RPC-memory specification problem, 
cf.~\cite{stolen1996using}.
 
Broy applied the \Focus-based theory to  introduce a logical basis for component-based system engineering \cite{Broy99alogical} as well as for 
specification of interface behaviour of multifunctional systems
\cite{Broy2010MSS}, i.e.,    systems that offer a variety of functions for different purposes and use cases. 
In \cite{Broy2010MSS}, service hierarchies specify multifunctional systems in terms of 
services (provided sub-functions) taking into account their mutual relationships and dependencies. Each service is specified independently and the specification is added to the service hierarchy, which then describes the functionality of multifunctional systems.

The Janus approach \cite{broy_janus,Broy2007services} was introduced 
to formally design of services and layered architectures, based 
 on the \Focus theory of distributed systems. 
 A Janus service, like a \Focus component, has a syntactic interface, but, 
 in comparison to a component, a service has a partial behaviour.

A methodology for modularised specification and verification of distributed time-triggered systems was proposed in \cite{botaschanjan2005towards,botaschanjan2006towards}. 
The authors applied \Focus to specify the systems formally. 

A number of \Focus-related works were focusing on refinement aspects:
from a general analysis of compositional refinement of interactive systems \cite{broy_refinement}  as well as the notions of abstraction and causality on the specification level  \cite{broy_time_abstr} to the refinement-based verification of interactive systems \cite{spichkova2008refinement,spichkova2009refinement,botaschanjan2008correctness} and its relation to the architectural aspects of software and system development \cite{spichkova2011refinement}. 
The corresponding methodology of architectural decomposition  was discussed in
\cite{spichkova2010architecture,spichkova2011decomp}. 
Transformation of semi-formal requirements to formal \Focus specifications  was proposed in \cite{spichkova2010semiformal}.

A specification and proof methodology  ``\Focus on Isabelle''
 \cite{spichkova2013we,spichkova,spichkova2008focus}. 
supports an alignment on the future proofs during specification phase
to make the proofs simpler and appropriate for application not only in theory but also in practice.  
Given a system represented in \Focus or \focust, the methodology
allows us to verify system properties by translating the specification to a Higher-Order Logic and 
subsequently using an interactive semi-automatic theorem prover~\cite{npw}. 
or the point of disagreement will be found. 
Another advantage of the methodology is a well-developed theory of composition, both for general 
and cryptographic properties \cite{spichkova2008formal,spichkova2012component}. 
 
Feilkas et al. presented  a methodology for the top-down 
development of automotive software systems \cite{feilkas2009top,feilkas2011refined}. 
The core artefact applied within the methodology was AutoFocus tool and the corresponding extension for test case generation. The methodology was elaborated using AutoFocus 2, however, it can be applied without any changes for AF3.
An extension of this \Focus-based methodology to the domain of cyber-physical systems was proposed in \cite{spichkova2012cyber}.

Vogelsang et al. \cite{vogelsang2014supporting} proposed a model-based approach
that starts from informal use cases and enables
a stepwise formalization of functional requirements 
to be linked to the architecture of the
system.  The approach utiled AF3 as the modelling framework.  

An approach introduced by Doby et al. \cite{dobi2015model} utilized \Focus to provide 
an efficient hazard and impact analysis for automotive mechatronics systems. 

Another approach based on \focust, allows analysis of component dependencies    \cite{spichkova2014formalisation}. This was later extended to framework for formal analysis of dependencies among services \cite{spichkova2014towards}.

The theory of processes extending the \Focus framework was introduced in 
\cite{spichkova2011focus}. 
This theory was further extended to have
a formal model of processes that is compatible with the component/ data flow view, cf. 
\cite{spichkova2015formal,mise2015process}. In many cases, it is beneficial to specify on the same abstraction level not only system components  
but also processes within the system, however, if we have to apply different frameworks to analyse both views,  
the system model becomes hard to read and to understand. 
The presented in \cite{mise2015process} approach provides 
a solution how to cover this gap and to  reconcile   component and process  views.

\section{\uppercase{Case Studies}}
\label{sec:applications}

The summaries of the first case studies in \Focus was presented approx. 25 years ago, in 1992 and 1994, cf. \cite{broy1992summary,broy1994summary}. 
The second summary was presented 5 years later, in 1997 \cite{broy1997summary}.

Specification and refinement of a buffer of length one was presented in \cite{broy1995buffer}.
The first version of the \Focus specification for a steam boiler system was introduced in \cite{broy1999boiler}.

A Trading System case study was used as common example 
for modelling approaches of component-based systems within the CoCoME contest to compare software component models. 
Broy at al. presented within this contest a systematic model-based approach of the engineering of distributed systems, which was based on the application of \Focus and AutoFocus2, cf. \cite{cocome}. The authors provide tool support around the AutoFocus2 tool,
that enables us to execute our specified models in a distributed environment
targeted to the CoCoME example.

Formal \Focus specifications of FlexRay and FTCom were introduced in \cite{kuhnel2007fault,kuhnel2006flexray,kuhnel2006upcoming}, where  
FlexRay is a time-triggered communication protocol,  and 
FTCom is the communication layer of a time-triggered 
operating system OSEKtime.
An operating system OSEKtime  
was developed by the European Automotive Consortium OSEK/VDX in accordance to the time-triggered paradigm. 
OSEK is a standards body, founded by German automotive company consortium,  
which included many industrial partners (e.g., BMW, Bosch, Siemens, etc.) as well as the University of Karlsruhe. 
The French automotive manufacturers Renault and PSA Peugeot Citroen had a similar consortium, VDX.
In 1994, a new consortium OSEK/VDX\footnote{\url{http://www.osek-vdx.org}} was created, based on OSEK and VDX.
FlexRay and FTCom were introduced for 
the fault-tolerant communication   by the FlexRay Consortium and
OSEK/VDX respectively.
Authors presented \Focus specifications of FlexRay and FTCom that allow us to argue about
the properties of FlexRay and FTCom in a formal manner. 
The \Focus specification of FlexRay was also verified methodology  ``\Focus on Isabelle'', discussed in Section \ref{sec:methodologies}, confirming that this 
specification of FlexRay fulfils the FlexRay requirements, cf. \cite{spichkova2006flexray}.
A number of case studies on the modelling of autonomous systems were presented in \cite{spichkova2015towards,spichkova2017autonomous,spichkova2016automotive}.

Within the methodology  ``\Focus on Isabelle'', three case studies were elaborated using the version of Isabelle framework of 2007:
\begin{itemize} 
\item data transmission (FlexRay communication protocol), 
\item process control (Steam Boiler System),  
\item memory and processing components (Automotive-Gateway System). 
\end{itemize}
These case studies were later conducted using the Isabelle  version of 2012, 
which allows to use the Isar language \cite{Isar} providing human-readable proofs in HOL), cf. \cite{FocusStreamsCaseStudies-AFP}. 
A further optimisation of the case studies on the verification level was proposed in \cite{spichkova2017human}: the authors introduced a human-oriented methodology for analysis of the dependencies between lemmas within the provided set of proofs, extending the ``\Focus on Isabelle'' methodology.

The case study introduced in \cite{CryptoBasedCompositionalProperties-AFP}, presents 
a \Focus formalisation of the security property of data secrecy along with  
the corresponding definitions and   Isabelle/HOL proofs.

B{\"o}hm et al. \cite{bohm2014formal} 
reported  successful results of a project conducted in collaboration between Siemens AG, fortiss GmbH and TU Munich. The goal was to evaluate SPES modeling framework (SPES MF)  implemented within AF3. B{\"o}hm et al. performed a case study, on modelling of requirements and functionality for a part of a Siemens train automation system. The results demonstrated advantages of application the SPES MF and AF3 within this context.

Campetelli et al. \cite{campetelli2015model} presented an industrial case study from the automation domain,
focusing on the control software components: 
a case example of a seawater desalination plant was
modelled in AF3 according to the proposed SPES development method.
 
Spichkova et al. \cite{spichkova2016formal} illustrated using formal models for intellingent speed validation and adaptation.
Formal specification of Chiminey platform \cite{Chiminey,yusuf2017chiminey}, which provides a reliable computing and data management service, as well as its refinements and extensions were presented in 
\cite{ChimineyICPADS,spichkova2016managing,spichkova2017towards}. 

Zamansky et. al. \cite{zamansky2016formal} reviewing some recent large-scale industrial projects in which formal methods (including \Focus and AutoFocus) have been successfully applied. The authors also covered some aspects of teaching formal methods for software engineering, including \Focus and AutoFocus, cf. \cite{spichkova2016teaching,simic2016enhancing}.

\section{\uppercase{Human-Oriented Aspects}}
\label{sec:hofm}

Sch{\"a}tz et al. \cite{schatz1996graphical} 
argued almost 20 years ago that that formal techniques are indeed useful for practical application, but they should be put to indirect use. 
To demonstrate this approach, the authors analysed   two pragmatic graphical description techniques, taken from the field of telecommunication.
 The analysis was  targeting on the information content of the techniques and their application in the process of specification development. The authors defined the techniques  formally, and introduced based on these formal definitions a number of  development steps and their graphical counterparts. This work can be seen as the first step towards graphical specification style within the \Focus framework as well as to the  AutoFocus tool.

An approach presented in \cite{hffm_spichkova,spichkova2013design}
aims to apply the engineering psychology achievements to the design of formal methods, focusing on the specification phase of a system development process.
Its core ideas originated from the analysis of the \Focus framework and also led to 
an extended version of the framework, \focust.  

\cite{spichkova2015human} introduced an research on incorporation of the human factors engineering into the software development process:  The authors proposed to apply the 
human factors analysis not only the level of requirements specification and formal system modelling, but also to guide various testing tasks.


\section{\uppercase{Conclusions}}
\label{sec:conclusions}

This paper provides a literature review of the methodological approaches, academic case studies and industrial applications of 
\Focus and \focust, 
the framework for formal specification and development of interactive systems, as well as
the AF3 (AutoFocus 3) modelling tool developed based on the  \Focus theory.

\Focus was introduced approx. 25 years ago, and extensively used since then in many academic and industrial projects.  
The literature review covers more than 80 publications on \Focus, \focust  and AutoFocus research, from 1992 when the first publication on \Focus appeared till 2017.

\bibliographystyle{abbrv} 

\vfill
 
\end{document}